My father is 5 times older than me. 12 years ago, he was 7 times older than me. How old is my father, and how old am I?
Answer by stanbon(75887)

My father is 5 times older than me. 12 years ago, he was 7 times older than me. How old is my father, and how old am I?
---
f = 5m
----
(f-12) = 7(m-12)
----
Substitute for "f" and solve for "m":
(5m-12) = 7m-84
2m = 72
m = 36 ("Me"'s age now)
---
f = 5m = 180 yrs old (Father's age now)
-----
I'm amazed.
---------------------
Cheers,
Stan H.